if the vant hoff factor of 0.01 m k2s04 solution is 2.64 the percentage of un dissociated k2so4 at this concentration is

Grade:12

1 Answers

Vikas TU
7 years ago
K2SO4    --------->               2K+           +    SO42-
Vant Hoff Factor =  2 + 1 = > 3
But 2.64 is theoretical value.
% of dissociated  = > 2.64/3 * 100 => 88%
Undissociated = > 100 – 88 = > 12%